OpenStudy (anonymous):

ok integral of INT/limits 1 to 0/ (e^x)^2

OpenStudy (anonymous):

rewrite?

OpenStudy (anonymous):

\[\int_0^1 e^{x^2}dx\]

OpenStudy (anonymous):

ah it is error function or something like that I can't do it

OpenStudy (agreene):

yeah, i think its like 1/2 sqrt(pi)*error(imaginary) or some such

OpenStudy (anonymous):

yep. that's what my calcu says.. 1.462651746 xD

OpenStudy (anonymous):

what is this now ?

OpenStudy (anonymous):

nonono, (e^x)(e^x) or (e^x)^2, integrate that from 1 to 0

OpenStudy (anonymous):

And I have a computer right in front of me as well as a ti-84+ so I dont want your calculator or wolfram alpha answers please :)

OpenStudy (agreene):

Oh, that is much simpler. (e^x)^2 = e^2x so: \[\int\limits_{0}^{1}e^{2x}dx=\frac{1}{2}e^2x\] take it to the limits: [1/2*e^(2)]-[1/2*e^(0)]=1/2e^2-1/2 factor it and you have: \[\frac{1}{2}(e^2-1)\]

OpenStudy (anonymous):

oh you sure that (e^x)^2=e^2x?

OpenStudy (agreene):

yes, that is one of the properties of exponents

OpenStudy (anonymous):

fsho